A new process for producing small precision parts is being studied. The process consists of mixing fine metal powder with a plastic binder, injecting the mixture into a mold, and then removing the binder with a solvent. These data are obtained on parts that should have a 1-inch diameter and whose standard deviation should not exceed 0.0025 inch. Perform all hypothesis tests at a 5% level of significance.

1.0030     0.9997     0.9990    1.0054     0.9991

1.0041     0.9988     1.0026     1.0032     0.9943

1.0021     1.0028     1.0002     0.9984     0.9999

What is the sample mean? (Round off to 5 decimal figures)

What is the sample standard deviation?

Formulate H1 for testing μ.

What is/are the critical region(s)?

What is the computed value of the test statistic?

Formulate H1 for testing for σ2

What is/are the critical region (s)?

What is the computed value of the test statistic?

                    What is the probability that the sample mean is at least 0.9988 if μ is known to be 1.0008  

                            and σ is unknown?

If σ is known to be 0.0022, find the probability that the sample standard deviation is at

                           most 0.0026?
Answer #1

parts. There is a sample of 15 The mean is given by : - Sym of all the sample values = 1.00304 1.0041+ 1.0021 +0.9997 +0.9988

t= 1.00084-1 D.00272/514 = (0.00084) x 3.741 0.00272 [t= 1.1553 ) for 14 defo t-tabulated value for 14 def ttab = 2.15 for 14

Tabulated value [x2 ARRRR = 23.685 > Mato at 5% level of significance accept Ho and a² f10.0025)? - » 7 Probability of sampleP(-505 < < I I = ploca <5.5) + Ploc Z<1) = 0.50+ 0.3413 = 0.8413